# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

### v1

@@ 6.1 Application of the Generalized Fermat-Torricelli Poblem
The dynamic sets in the absence of wind, corresponding to the UAV’s nominal speed $r$ (m/s), are represented by the set:
@@
Taking wind into account, with wind vector $\mathcal{S}\in\mathbb{R}^{2}$ and $\mathcal{S}=(s_{1},s_{2})$, the effective dynamic set—representing the UAV’s velocity under wind conditions—is given by
@@
Figure 2: Case 1 in Table 1: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic set with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAV under windy conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction
@@
Dispatching the UAV from the mobile station involves movement in the opposite direction, so the corresponding dynamic set is given by $F_{i}$ for $i=1,2,3$, as stated in Proposition 8.9. We assume that the wind vector satisfies the condition $\left\|\mathcal{S}\right\|_{2}<r$, ensuring that the UAV can still operate effectively under the given windy condition. Therefore, the generalized Fermat-Torricelli problem is represented as (for simplicity, we assume $\bar{F}=-F$):
$\min_{x\in\Omega_{0}}T(x)\ \ \textnormal{where}\ \ T(x)=\sum_{i=1}^{3}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\rho_{-F_{i}}^{\Omega_{i}}(x)=\sum_{i=1}^{3}\rho_{F}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\rho_{\bar{F}}^{\Omega_{i}}(x),$ (6.2)
where $x=(x_{1},x_{2})\in\mathbb{R}^{2}$. To solve the problem (6.2), we divide the solution approach into several steps.
Step 1: Obtaining the Minkowski gauge function $\rho_{F}(x):$
We begin by computing the Minkowski gauge function (2.5). The function $\rho_{F}(x)$ is obtained as the optimal value of the following optimization problem:
$\begin{split}&\min\ t\\ \text{s.t.}\quad&(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}=r^{2},\\ &x_{1}=tf_{1},\\ &x_{2}=tf_{2},\\ &t\geq 0.\end{split}$ (6.3)
According to Proposition 8.8, the first constraint in (6.3) enforces that the velocity vector $(f_{1},f_{2})$ lies on the boundary of the dynamic set $F$. To derive a closed-form expression for $\rho_{F}(x)$, we can substitute $f_{1}=x_{1}/t$ and $f_{2}=x_{2}/t$ into the first constraint in (6.3), yielding:
$\begin{split}&\ (\frac{x_{1}}{t}-s_{1})^{2}+(\frac{x_{2}}{t}-s_{2})^{2}=r^{2}\\ \Rightarrow&\ (x_{1}-ts_{1})^{2}+(x_{2}-ts_{2})^{2}=r^{2}t^{2}\\ \Rightarrow&\ t^{2}(s_{1}^{2}+s_{2}^{2}-r^{2})-2(s_{1}x_{1}+s_{2}x_{2})t+(x_{1}^{2}+x_{2}^{2})=0\end{split}$ (6.4)
By solving the quadratic equation obtained from (6.4), we have:
$\rho_{F}(x)=t=\frac{(s_{1}x_{1}+s_{2}x_{2})-\sqrt{x_{1}^{2}(r^{2}-s_{2}^{2})+x_{2}^{2}(r^{2}-s_{1}^{2})+2s_{1}s_{2}x_{1}x_{2}}}{s_{1}^{2}+s_{2}^{2}-r^{2}}$ (6.5)
The value of $t$ obtained from (6.5) is the optimal solution to the problem in (6.3), and is obviously nonnegative.
Step 2: Obtaining the subdiferential of $\rho_{F}(x)$ and $\rho_{F}^{\Omega}(x)$ functions:
In this step, we compute the subdifferentials of the functions $\rho_{F}(x)$ and $\rho_{F}^{\Omega}(x)$, which are necessary for implementing the subgradient algorithm (Algorithm 1)—a well-known method in convex optimization—in order to find the minimizer of the problem in (6.2). Let $\alpha=\frac{1}{s_{1}^{2}+s_{2}^{2}-r^{2}}$, and define
$\beta(x_{1},x_{2})=x_{1}^{2}(r^{2}-s_{2}^{2})+x_{2}^{2}(r^{2}-s_{1}^{2})+2s_{1}s_{2}x_{1}x_{2}.$
Then, for $x\neq 0$, the gradient of the function $\rho_{F}(x)$ in (6.5) is given by:
@@
Therefore, the subdifferential of the function $\rho_{F}^{\Omega_{i}}$ can be expressed as follows:
$\partial\rho_{F}^{\Omega_{i}}(x)=\begin{cases}N(x;\Omega_{i})\cap\mathcal{V}&\text{if }x\in\Omega_{i},\\ \{\nabla\rho_{F}(x-\omega_{i})\}&\text{if }x\notin\Omega_{i},\end{cases}\quad\text{for }i=1,2,3,$ (6.7)
where $\omega_{i}\in\Pi_{F}(x;\Omega_{i})$, and $\mathcal{V}$ is defined in Proposition 3.4. Note that in (6.7), for $x\notin\Omega_{i}$, $x-\omega_{i}\neq 0$ and the subdifferential is obtained using Proposition 3.6. For $x\in\Omega_{i}$, we have $x=\Pi_{F}(x;\Omega_{i})$, and (6.6) does not provide the gradient at the origin. Therefore, we apply Proposition 3.4 to compute the subdifferential of $\rho_{F}^{\Omega_{i}}(x)$ at $x\in\Omega_{i}$.
@@
Step 3: Obtaining the optimal solution by subgradient algorithm:
To determine the optimal location of the mobile station, we apply the subgradient algorithm (Algorithm 1) to find the minimizer of the generalized Fermat–Torricelli problem (6.2). According to the structure of the subgradient algorithm, we need to compute a subgradient of $T(x)$. To do so, we calculate the subdifferential of $T(x)$ by applying the subdifferential sum rule (Theorem 2.4), and we obtain:
Therefore, at each iteration $k$ of the subgradient algorithm, a subgradient $v^{k}\in\partial T(x^{k})$ is computed as
$v^{k}=\sum_{i=1}^{3}v_{i}^{k}+\sum_{i=1}^{3}\bar{v}_{i}^{k},$ (6.9)
where each $v_{i}^{k}\in\partial\rho_{F}^{\Omega_{i}}(x^{k})$ and $\bar{v}_{i}^{k}\in\partial\rho_{\bar{F}}^{\Omega_{i}}(x^{k})$ is determined based on the formula in (6.7) (taking into account that $0\in N(x^{k};\Omega_{i})\cap\mathcal{V}$):
$v_{i}^{k}=\begin{cases}0&\text{if }x^{k}\in\Omega_{i},\\ \nabla\rho_{F}(x^{k}-\omega_{i}^{k})&\text{if }x^{k}\notin\Omega_{i},\end{cases}\quad\text{for }i=1,2,3,$ (6.10)
with $\omega_{i}^{k}\in\Pi_{F}(x^{k};\Omega_{i})$. Also, we have:
$\bar{v}_{i}^{k}=\begin{cases}0&\text{if }x^{k}\in\Omega_{i},\\ \nabla\rho_{\bar{F}}(x^{k}-\bar{\omega}_{i}^{k})&\text{if }x^{k}\notin\Omega_{i},\end{cases}\quad\text{for }i=1,2,3,$ (6.11)
where $\bar{\omega}_{i}^{k}\in\Pi_{\bar{F}}(x^{k};\Omega_{i})$. Note that by solving the problem in (6.3) with $F$ replaced by $\bar{F}$, we can compute both $\rho_{\bar{F}}(x)$ and its gradient $\nabla\rho_{\bar{F}}(x)$.
To illustrate the procedure, we focus on Case 1 in Table 1 and evaluate $v^{k}\in\partial T(x^{k})$ for $k=1$, where the dynamic sets are specified by Info 1 in Table 4 and the wind vector is given by $\mathcal{S}=(-0.6,0.6)$. Therefore, the dynamic sets for this UAV, corresponding to back-and-forth movements, are given by:
$\begin{split}F=\{(f_{1},f_{2})\in\mathbb{R}^{2}|(f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 1\},\\ \bar{F}=\{(f_{1},f_{2})\in\mathbb{R}^{2}|(f_{1}-0.6)^{2}+(f_{2}+0.6)^{2}\leq 1\}.\end{split}$ (6.12)
Let the initial point be $x^{1}=(100,100)$, and consider the computation of $v_{1}^{1}$. Since $x^{1}\notin\Omega_{1}$, the formula in (6.10) implies that it is sufficient to compute $\nabla\rho_{F}(x^{1}-\omega_{1}^{1})$. With $\omega_{1}^{1}=(45,335)$, we have
$x^{1}-\omega_{1}^{1}=(100-45,\,100-335)=(55,-235).$
Applying the formula in (6.6), we obtain
Using the same procedure, we obtain:
[…]
Finally, $v^{1}$ can be computed using (6.9), yielding
$v^{1}=(-6.66,1.43).$
Following this procedure, we compute $v^{k}$ at each iteration. By choosing $\alpha^{k}=\tfrac{1}{k}$ and applying Algorithm 1, we obtain an approximate solution to the generalized Fermat-Torricelli problem (6.2).
For Case 1 in Table 1, we first simplify the problem by neglecting the wind effect and considering the dynamic sets for the back-and-forth movement as $F=\bar{F}=F_{0}$, as illustrated in Fig. 2(b). In this scenario, the approximate solution is $(252,117)$, shown in Fig. 2(a). If this solution is applied in a real-world scenario with wind, the resulting transition time is 4039 seconds. When the wind effect is incorporated, the dynamic sets are defined as in (6.12), illustrated in Fig. 2(d). The optimal solution then shifts to $(242,52)$, depicted in Fig. 2(c), reducing the objective value to 3960 seconds. It should be noted that, since the sets $\Omega_{i}$ for $i=1,2,3$ are not strictly convex, the uniqueness of the problem’s solution cannot be guaranteed, as stated in Proposition 8.18.
Table 1 presents ten cases with varying settings and parameters to evaluate the efficiency of the generalized Fermat–Torricelli problem in accounting for wind effects. For simplicity, all cases are conducted on a small scale, with the UAV’s nominal speed set to $1\text{(m/s)}$; these parameters can be appropriately scaled to reflect the real scenarios. The results indicate that, for single-UAV operations, the generalized Fermat–Torricelli problem reduces the transition time by approximately $3.3\%$.
Algorithm 1 subgradient Algorithm
@@
6.2 Application of the Generalized Sylvester Problem
In the context of Section 6.1, we consider a scenario in which the rescue team deploys a heterogeneous fleet of three UAVs. These UAVs are launched at the earliest possible time following a disaster, and the objective is to determine the optimal location for a mobile station (truck) to collect them as quickly as possible. Timely retrieval of the UAVs enables rapid access to critical data, which is essential for effective response and decision-making in emergency situations. We consider this problem by the following setting (corresponding to Case 2 in Table 2):
@@ 6.2 Application of the Generalized Sylvester Problem
$\Omega_{1}$ $=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-30)^{2}}{10^{2}}+\frac{(y-350)^{2}}{10^{2}}\leq 1\},$
$\Omega_{2}$ $=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-210)^{2}}{10^{2}}+\frac{(y-10)^{2}}{10^{2}}\leq 1\},$
$\Omega_{3}$ $=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-550)^{2}}{10^{2}}+\frac{(y-200)^{2}}{10^{2}}\leq 1\}.$
For dynamic sets, we have:
$\ F_{1}$ $=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 2^{2}\},$
$\ F_{2}$ $=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 1^{2}\},$
$\ F_{3}$ $=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 3^{2}\}.$
We set 1(m/s) as the nominal speed of the slowest UAV (which is represented by $F_{2}$); consequently, the rescue team is equipped with two additional UAVs that are twice and three times faster, respectively. In addition, we consider the wind vector as $\mathcal{S}=(-0.6,0.6)$.
To determine the optimal location of the mobile station, we formulate the problem as the following generalized Sylvester problem, which is a variant of the generalized SFT problem:
$\min_{x\in\Omega_{0}}Y(x)\ \ \ \textnormal{where}\ \ \ Y(x):=\max\{\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ \ i=1,2,3\}.$ (6.13)
To apply Algorithm 1 for approximating a minimizer of the problem in (6.13), we need to compute a subgradient of $Y(x)$ at each iteration $k$. This is done by following the procedure described in Section 6.1 to compute $v_{i}^{k}\in\partial\rho_{F_{i}}^{\Omega_{i}}(x^{k})$. Subsequently, $v^{k}\in\partial Y(x^{k})$ is obtained using the formula derived from Theorem 2.3:
@@
Initially, we consider the case where the wind effect is ignored ($\mathcal{S}=0$) and the dynamic sets are illustrated in Fig. 3(b). The approximated solution in this case is $(175,138)$, as shown in Fig. 3(a), which corresponds to a transition time of $212$ seconds in the real scenario under windy conditions. When the wind effect is taken into account with $\mathcal{S}=(-0.6,0.6)$, the dynamic sets are modified as depicted in Fig. 3(d), yielding the approximated optimal location $(108,216)$, as illustrated in Fig. 3(c). The transition time for $(108,216)$ is $124$ seconds, representing a $41.5\%$ reduction in transition time. This demonstrates the effectiveness of the generalized Sylvester problem in addressing this challenge.
To evaluate the performance of the generalized Sylvester problem in mitigating wind effects, 14 cases with varying settings and parameters are considered, as summarized in Table 2. For simplicity, all cases are performed on a small scale, assigning 11(m/s) to the slowest UAV’s nominal speed; these values can be scaled to represent real-world scenarios. The results demonstrate that the generalized Sylvester problem significantly reduces transition time, achieving an approximate $84\%$ improvement. In Cases 1 to 7 of Table 2, all parameters are identical except for the length of the wind vector. The results of these cases are illustrated in Fig. 5(b), which highlights the impact of wind speed. As expected, at higher wind speeds, applying the generalized Sylvester problem resulted in greater efficiency and a larger reduction in transition time.
6.3 Application of the Generalized SFT Problem

### v2

@@ 6.1 Applications of the Generalized Fermat-Torricelli Problem
The dynamic sets in the absence of wind, corresponding to the UAV’s nominal speed $r$ (m/s), are represented by the ball of the plane:
@@
Taking wind into account, with wind vector $\mathcal{S}=(s_{1},s_{2})\in\mathbb{R}^{2}$, the effective dynamic set–representing the UAV’s velocity under wind conditions–is given by
@@
Figure 3: Case 1 in Table 1: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic set with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAV under wind conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction
@@
Dispatching the UAV from the mobile station involves movement in the opposite direction, so the corresponding dynamic set is given by $F_{i}$ for $i=1,2,3$, as stated in Proposition 8.8 from the Appendix. We assume that the wind vector satisfies the condition $\left\|\mathcal{S}\right\|_{2}<r$, ensuring that the UAV can still operate effectively under the given wind condition. Therefore, the generalized Fermat-Torricelli problem is represented as (assuming for simplicity that $\bar{F}=-F$):
$\min_{x\in\Omega_{0}}T(x)\ \ \textnormal{with}\ \ T(x)=\sum_{i=1}^{3}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\rho_{-F_{i}}^{\Omega_{i}}(x)=\sum_{i=1}^{3}\rho_{F}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\rho_{\bar{F}}^{\Omega_{i}}(x),$ (6.2)
where $x=(x_{1},x_{2})\in\mathbb{R}^{2}$. To solve (6.2), we split the solution procedure into several steps.
Step 1: Computing the Minkowski gauge function $\rho_{F}(x).$
The classical Minkowski gauge $\rho_{F}(x)$ in (2.4) is the optimal value of the following problem:
$\begin{split}&\min\ t\;\mbox{ subject to}\\ &(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}=r^{2},\\ &x_{1}=tf_{1},\;x_{2}=tf_{2},\;t\geq 0.\end{split}$ (6.3)
According to Proposition 8.7, the first constraint in (6.3) enforces that the velocity vector $(f_{1},f_{2})$ lies on the boundary of the dynamic set $F$. To derive a closed-form expression for $\rho_{F}(x)$, we substitute $f_{1}=x_{1}/t$ and $f_{2}=x_{2}/t$ into the first constraint in (6.3) and get
$\begin{split}&\ \displaystyle\Big(\frac{x_{1}}{t}-s_{1}\Big)^{2}+\Big(\frac{x_{2}}{t}-s_{2}\Big)^{2}=r^{2}\\ \Longrightarrow&\ (x_{1}-ts_{1})^{2}+(x_{2}-ts_{2})^{2}=r^{2}t^{2}\\ \Longrightarrow&\ t^{2}(s_{1}^{2}+s_{2}^{2}-r^{2})-2(s_{1}x_{1}+s_{2}x_{2})t+(x_{1}^{2}+x_{2}^{2})=0.\end{split}$ (6.4)
Solving the quadratic equation from (6.4) gives us
$\rho_{F}(x)=t=\displaystyle\frac{(s_{1}x_{1}+s_{2}x_{2})-\sqrt{x_{1}^{2}(r^{2}-s_{2}^{2})+x_{2}^{2}(r^{2}-s_{1}^{2})+2s_{1}s_{2}x_{1}x_{2}}}{s_{1}^{2}+s_{2}^{2}-r^{2}}.$ (6.5)
The value of $t$ obtained from (6.5) is the optimal solution to the problem in (6.3), and it is obviously nonnegative.
Step 2: Computing subgradients of the functions $\rho_{F}(x)$ and $\rho_{F}^{\Omega}(x)$.
Here we compute subgradients of the functions $\rho_{F}(x)$ and $\rho_{F}^{\Omega}(x)$, which are used below in the subgradient algorithm (Algorithm 1) to solve problem (6.2). Define $\alpha:=\frac{1}{s_{1}^{2}+s_{2}^{2}-r^{2}}$ and
$\beta(x_{1},x_{2}):=x_{1}^{2}(r^{2}-s_{2}^{2})+x_{2}^{2}(r^{2}-s_{1}^{2})+2s_{1}s_{2}x_{1}x_{2}.$
Then for any $x\neq 0$, the gradient of the function $\rho_{F}(x)$ in (6.5) is calculated by
@@
Therefore, the subdifferential of the function $\rho_{F}^{\Omega_{i}}$ are calculated by Theorems 3.4 and Proposition 3.6 as follows:
$\partial\rho_{F}^{\Omega_{i}}(x)=\begin{cases}N(x;\Omega_{i})\cap\mathcal{V}&\text{if }x\in\Omega_{i},\\ \{\nabla\rho_{F}(x-\omega_{i})\}&\text{if }x\notin\Omega_{i},\end{cases}\quad i=1,2,3,$ (6.7)
where $\omega_{i}\in\Pi_{F}(x;\Omega_{i})$ and $x-\omega_{i}\neq 0$ if $x\notin\Omega_{i}$.
@@
Step 3: Computing the optimal solution by subgradient algorithm.
To determine the optimal location of the mobile station, we apply the subgradient algorithm (see Algorithm 1 below) allowing us to find an optimal solution to the generalized Fermat–Torricelli problem under consideration in (6.2).
Algorithm 1 Subgradient Method
@@
According to the structure of the subgradient algorithm, we need to compute a subgradient of $T(x)$. To do so, we calculate the subdifferential of $T(x)$ by applying the subdifferential sum rule from Theorem 2.4, which gives us the representation
Therefore, at each iteration $k$ of the subgradient algorithm, a subgradient $v^{k}\in\partial T(x^{k})$ can be calculated as follows:
$v^{k}=\sum_{i=1}^{3}v_{i}^{k}+\sum_{i=1}^{3}\bar{v}_{i}^{k},$ (6.9)
where $v_{i}^{k}\in\partial\rho_{F}^{\Omega_{i}}(x^{k})$ and $\bar{v}_{i}^{k}\in\partial\rho_{\bar{F}}^{\Omega_{i}}(x^{k})$ are taken from (6.7) as
$v_{i}^{k}=\begin{cases}0&\text{if }x^{k}\in\Omega_{i},\\ \nabla\rho_{F}(x^{k}-\omega_{i}^{k})&\text{if }x^{k}\notin\Omega_{i},\end{cases}\quad i=1,2,3,$ (6.10)
with $\omega_{i}^{k}\in\Pi_{F}(x^{k};\Omega_{i})$. Similarly we get
$\bar{v}_{i}^{k}=\begin{cases}0&\text{if }x^{k}\in\Omega_{i},\\ \nabla\rho_{\overline{F}}(x^{k}-\bar{\omega}_{i}^{k})&\text{if }x^{k}\notin\Omega_{i},\end{cases}\quad i=1,2,3,$ (6.11)
where $\bar{\omega}_{i}^{k}\in\Pi_{\overline{F}}(x^{k};\Omega_{i})$. Note that by solving the problem in (6.3) with $F$ replaced by $\overline{F}$, we can compute both $\rho_{\overline{F}}(x)$ and its gradient $\nabla\rho_{\overline{F}}(x)$.
To illustrate the computation procedure, we focus for definiteness on Case 1 in Table 1 and evaluate $v^{k}\in\partial T(x^{k})$ for $k=1$, where the dynamic sets are specified by Info 1 in Table 4 and the wind vector is given by $\mathcal{S}=(-0.6,0.6)$. Therefore, the dynamic sets for this UAV, corresponding to back-and-forth movements, are given by
$\begin{split}F=\big\{(f_{1},f_{2})\in\mathbb{R}^{2}\big|\;(f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 1\big\},\\ \bar{F}=\big\{(f_{1},f_{2})\in\mathbb{R}^{2}\;\big|\;(f_{1}-0.6)^{2}+(f_{2}+0.6)^{2}\leq 1\big\}.\end{split}$ (6.12)
Starting with $x^{1}=(100,100)\notin\Omega_{1}$, compute $v_{1}^{1}$ by (6.10) With $\omega_{1}^{1}=(45,335)$. This gives us
$x^{1}-\omega_{1}^{1}=(100-45,\,100-335)=(55,-235).$
Then it follows from (6.6) that
Using the same procedure, we get
[…]
Finally, $v^{1}$ can be computed by using (6.9), which brings us to
$v^{1}=(-6.66,1.43).$
In the same way, we compute $v^{k}$ at each iteration. Then Algorithm 1 with the standard choice of $\alpha^{k}=\tfrac{1}{k}$ allows us to solve the generalized Fermat-Torricelli problem (6.2).
To proceed with Case 1 in Table 1, we first simplify the problem by neglecting the wind effect and considering the dynamic sets for the back-and-forth movement as $F=\bar{F}=F_{0}$ as illustrated in Fig. 3(b). In this scenario, the approximate solution is $(252,117)$ as shown in Fig. 3(a). If this solution is applied in a real-world scenario with wind, the resulting transition time is 4039 seconds. When the wind effect is incorporated, the dynamic sets are defined as in (6.12); see Fig. 3(d). The optimal solution then shifts to $(242,52)$, depicted in Fig. 3(c), reducing the objective value to 3960 seconds. Note that, since the sets $\Omega_{i}$ for $i=1,2,3$ are not strictly convex, the uniqueness of solutions to this problem cannot be generally guaranteed.
Table 1 presents ten cases with varying settings and parameters to evaluate the efficiency of the generalized Fermat–Torricelli problem in accounting for wind effects. For simplicity, all the cases are conducted on a small scale, with the UAV’s nominal speed set to $1\text{(m/s)}$; these parameters can be appropriately scaled to reflect the real scenarios. The results indicate that, for single-UAV operations, the generalized Fermat–Torricelli problem reduces the transition time by approximately $3.3\%$.
6.2 Applications of the Generalized Sylvester Problem
In the context of Section 6.1, we consider a scenario in which the rescue team deploys a heterogeneous fleet of three UAVs. These UAVs are launched at the earliest possible time after a disaster, and the goal is to determine an optimal location for a mobile station to collect them as quickly as possible. Timely retrieval of the UAVs enables rapid access to critical data, which is essential for effective response and decision-making in emergency situations. We consider this problem in the following setting corresponding to Case 2 from Table 2:
@@ 6.2 Applications of the Generalized Sylvester Problem
$\Omega_{1}$ $=\Big\{(x,y)\in\mathbb{R}^{2}\ \Big|\ \frac{(x-30)^{2}}{10^{2}}+\frac{(y-350)^{2}}{10^{2}}\leq 1\Big\},$
$\Omega_{2}$ $=\Big\{(x,y)\in\mathbb{R}^{2}\ \Big|\ \frac{(x-210)^{2}}{10^{2}}+\frac{(y-10)^{2}}{10^{2}}\leq 1\Big\},$
$\Omega_{3}$ $=\Big\{(x,y)\in\mathbb{R}^{2}\ \Big|\ \frac{(x-550)^{2}}{10^{2}}+\frac{(y-200)^{2}}{10^{2}}\leq 1\Bigg\}.$
For the dynamic sets, we have the following:
$\ F_{1}$ $=\big\{(f_{1},f_{2})\in\mathbb{R}^{2}\ \big|\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 2^{2}\big\},$
$\ F_{2}$ $=\big\{(f_{1},f_{2})\in\mathbb{R}^{2}\ \big|\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 1^{2}\big\},$
$\ F_{3}$ $=\big\{(f_{1},f_{2})\in\mathbb{R}^{2}\ \big|\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 3^{2}\big\}.$
Let us set 1(m/s) as the nominal speed of the slowest UAV, which is represented by $F_{2}$. Consequently, the rescue team is equipped with two additional UAVs that are twice and three times faster, respectively. In addition, consider the wind vector given by $\mathcal{S}=(-0.6,0.6)$.
To find an optimal location of the UAV station, we formulate this model as the following generalized Sylvester problem, which is a variant of the generalized SFT problem (2.6):
$\min_{x\in\Omega_{0}}Y(x)\ \ \ \textnormal{with}\ \ \ Y(x):=\max\big\{\rho_{F_{i}}^{\Omega_{i}}(x)\ \big|\ \ i=1,2,3\big\}.$ (6.13)
Then Algorithm 1 allows us to determine an approximate minimizer of (6.13). To proceed, we need to compute a subgradient of $Y(x)$ at each iteration $k$. Following the procedure described in Section 6.1 gives us the calculation of $v_{i}^{k}\in\partial\rho_{F_{i}}^{\Omega_{i}}(x^{k})$. Subsequently, $v^{k}\in\partial Y(x^{k})$ is obtained by using the formula derived from Theorem 2.3:
@@
Initially, we consider the case where the wind effect is ignored ($\mathcal{S}=0$) and the dynamic sets are illustrated in Fig. 4(b). The approximated solution in this case is $(175,138)$, as shown in Fig. 4(a), which corresponds to the transition time of $212$ seconds in the real scenario under windy conditions. When the wind effect is taken into account with $\mathcal{S}=(-0.6,0.6)$, the dynamic sets are modified as depicted in Fig. 4(d), which gives us the approximated optimal location $(108,216)$ as illustrated in Fig. 4(c). The transition time for $(108,216)$ is $124$ seconds, representing a $41.5\%$ reduction in transition time. This clearly demonstrates the effectiveness of the generalized Sylvester problem in addressing this challenge.
To evaluate the performance of the generalized Sylvester problem in mitigating wind effects, 14 cases with varying settings and parameters are considered as summarized in Table 2. For simplicity, all cases are performed on a small scale, assigning 11(m/s) to the slowest UAV nominal speed; these values can be scaled to represent real-world scenarios. The results demonstrate that the generalized Sylvester problem significantly reduces transition time, achieving an approximate $84\%$ improvement. In Cases 1–7 of Table 2, all parameters are identical except for the length of the wind vector. The results of these cases are illustrated in Fig. 6(b), which highlights the impact of wind speed. As expected, applying the generalized Sylvester model at higher wind speeds resulted in greater efficiency and a larger reduction in transition time.
6.3 Applications of the Generalized SFT Problem